OpenStudy (anonymous):

Which of the following is the radical expression of a to the four ninths power ? 4a9 9a4 fourth root of a to the ninth power ninth root of a to the fourth power

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

x^(a/b) = (x^a)^(1/b)

OpenStudy (whpalmer4):

for example, 4^(3/2) = the square root of 4^3

OpenStudy (mertsj):

Just remember: top number is the power, bottom number is the root

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

it is the second one DI_KAZ

OpenStudy (whpalmer4):

@michaelaepalmore might want to read the discussion again...

OpenStudy (mertsj):

So that means that 4 is the power and 9 is the root.

OpenStudy (anonymous):

IS it a?

OpenStudy (whpalmer4):

top number is power, bottom number is root. a to the 4/9 power means the 9th root of a to the 4th power Is that answer A?

OpenStudy (anonymous):

that is d

OpenStudy (whpalmer4):

answer A is likely to be 4a^9, I suspect, which is 4 times a to the 9th power. a totally different beast!

OpenStudy (whpalmer4):

is the 9 raised up higher than the a wherever the problem is appearing for you? if so, that's an exponent, and the convention is to represent it with an ^